OpenStudy (anonymous):

One thousand dollars is invested at 12% interest compounded annually. Determine how much investment is worth after 10 years.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\Huge\color{magenta }{ \sf A=P(1+ \frac{r}{n})^{n\times t} }\) A = amount of money accumulated after n years, including interest. P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. n = number of times the interest is compounded per year \(\Huge\color{blue}{ \sf In~~~your~~case, }\) \(\large\color{blue}{ \sf P = 1000 }\) \(\large\color{blue}{ \sf r = 0.12 }\) \(\large\color{blue}{ \sf t=10 }\) \(\large\color{blue}{ \sf n=1 }\) plug in the numbers into the formula (above in pink) and solve for A (the final amount of money after 10 years)

OpenStudy (anonymous):

I get \[A=1000(1+\frac{ 0.12 }{ 1 })^{10}\]

OpenStudy (anonymous):

and that equals $23,883.87?

OpenStudy (solomonzelman):

 https://www.google.com/#q=(1.12)%5E10 and then times 1000